OpenStudy (anonymous):

What is the equation of a line that passes through points (-1, 2) and (5, 2)? y - 2 = 0 x - 2 = 0 y - x - 2 = 0

jigglypuff314 (jigglypuff314):

first get what was given to you into an equation that you know...

jigglypuff314 (jigglypuff314):

so you would first have to find the slope of the line. \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]when given the points (x1, y1) and (x2, y2) so when given the points (-1, 2) and (5, 2) you would plug in x1 = -1 , y1 = 2 , x2 = 5 , y2 = 2 but when you try to get the slope, you get a 0 in the numerator therefore you're slope would be 0 (a horizontal line) (note that if you get a 0 in the denominator, you would get a vertical line) then slope-intercept form is: y = mx + b when m = slope and b = y-intercept but slope = 0 so y = 0 + b y = b so you would just plug in the y value of a given coordinate for b (since it is a horizontal line) (-1, 2) ^ so y = 2 then to match it up with your options get all values to one side so minus 2 from both sides... y = 2 - 2 -2 y -